G-081 field season 2010-2011

Logbook entries (add new)

Date Title Preview People
April 8, 2011 importing tinytag data Haven't checked for actual logbook entry yet, just importing the file.   edit
April 8, 2011 importing tinytag data Need to check written notes for details.   edit
April 8, 2011 importing tinytag data Need to check in written logs for logbook entry.   edit
April 8, 2011 importing tinytag data Need to enter paper record.   edit
Dec. 31, 2010 Warren cleanup, condensate sampling, etc - Added some things to RMY to campbell datalogger in Warren: RMY anemometer, 107... Aaron Curtis, Nelia Dunbar, edit
Dec. 28, 2010 Campbell and DTS - Put 5 medium batteries down Warren for Campbell with Nial, replaced DTS batts,... Aaron Curtis, Yves Mousallam, edit
Dec. 27, 2010 Sauna Cave bio sampling and tinytag install Rigged Sauna cave. Used rope from Kachina, three screws, bollard on ice block be... Nial Peters, Aaron Curtis, edit
Dec. 19, 2010 Tinytag replacement - Re-installed tinytags in Kachina V1 and Kachina Big Crack. Nial Peters, Aaron Curtis, edit
Dec. 19, 2010 Failed to recover mouse cave tinytag Tried to find tinytag in Mouse Cave. Cave has shrunk. Got in but couldn't find t... Nial Peters, Aaron Curtis, edit
Dec. 18, 2010 DTS download, datalogger start - Down Warren - Downloaded DTS - Started dataloggers with own laptop - Moved ... Aaron Curtis, Kayla Iacovino, edit
Dec. 18, 2010 Tourist trip for renewable energy people Introduced renewable energy people to Hut Cave (names?). Aaron Curtis, edit
Dec. 17, 2010 Warren gas sampling and attempted datalogger start - Gas sampling at both vents - Tried to start dataloggers with HP laptop but ba... Aaron Curtis, Nial Peters, edit
Dec. 16, 2010 Datalogger replacement Replaced dataloggers. Matt Zimmerer, Aaron Curtis, edit
Dec. 15, 2010 DTS cable installation in Warren Cave One of the trips to install a fiber optic distributed temperature sensing system... Nelia Dunbar, Bill McIntosh, Aaron Curtis, Matt Zimmerer, edit
Dec. 10, 2010 Ice tower ridge trip Re-deployed logger (UTC) Aaron Curtis, edit
Dec. 10, 2010 Field notes 10 Dec Replaced solar controller, replaced and started tinytag. Nial Peters, Aaron Curtis, edit
Dec. 10, 2010 Field notes 10 Dec: Kachina Nausea power dead, so Nial came with me to ITR and Kachina. Found Tinytag in mai... Nial Peters, Aaron Curtis, edit
Dec. 9, 2010 HOBO removal and gas sampling in hut cave 4x gas samples from main vent: G001 through G004 HOBO - CH1 0121117, at V2 - ...   edit
Jan. 2, 2010 Kachina wrap-up - Down Kachina with Nial and Nels. - Set up condensate sampling in V1 - Downlo... Nial Peters, Nels Iverson, Aaron Curtis, edit
Jan. 1, 2010 Measurement of temp profile into wall - Bad weather in the morning. - Afternoon: setup crater cam on MEVO site - Aft... Aaron Curtis, edit
loading page...